OpenStudy (anonymous):

in how many ways can 5 different trees be planted in a circle

OpenStudy (kropot72):

The information here will enable you to solve it: http://tutors4you.com/circularpermutations.htm

OpenStudy (ikram002p):



OpenStudy (ikram002p):

:D so its (5-1)! i guess

ganeshie8 (ganeshie8):

lol i was thinking you're going to explain :o

OpenStudy (ikram002p):

hehe well yeah wanna explain :P

OpenStudy (uri):

There is a formula for the circle (n-1)! (4)! =4*3*2*1=

OpenStudy (ikram002p):

:D

ganeshie8 (ganeshie8):

that link has nice explanation : number of circular arrangements of \(n\) different objects = \(\dfrac{n!}{n}\) because for each circular arrangement, there will be \(n\) different linear arrangements

OpenStudy (ikram002p):

yep ,cool